ย Consider any two positive real numbers and call it ๐‘Ž and ๐‘. Then consider the function defined asย ย 

๐‘“(๐‘ฅ) = {

๐‘Ž, 0 โ‰ค ๐‘ฅ < 1 ๐‘,ย ย ย ย ย ย ย ย ย ย ๐‘ฅ = 1 Find the ๐‘ˆ(๐‘ƒ,๐‘“) and ๐ฟ(๐‘ƒ,๐‘“) for the partition ๐‘ƒ = {๐‘ฅ0,๐‘ฅ1,โ€ฆ๐‘ฅ๐‘›} of [0,1]. Also check whether the function is Riemann integrable over [0,1] or not.ย 


1
Expert's answer
2021-07-16T16:04:35-0400

"Let \space P=\{0, \frac{1}{n},\frac{2}{n},...,\frac{r-1}{n}, \frac{r}{n}...,\frac{n}{n}=1\}" be any partition of [0,1]

The rth sub interval of [0,1] "= I_r=[\frac{r-1}{n},\frac{r}{n}] ; r= 1,2,...n"

If "\delta_r" be the length of Ir then, "\delta = \frac{r}{n}- \frac{r-1}{n}= \frac{1}{n}"

Let Mr and mr be respectively the supremum and the infimum of f in Ir . Since f is increasing on [0,1] we have

"M_r = supf(x) =\frac{r^3}{n^3}\\
m_r = inff(x) =\frac{(r-1)^3}{n^3}\\
U(f,p)= \sum^n_{r=1}M_rdr=\sum_{r=1}^n \frac{r^3}{n^3}\frac{1}{n}=\frac{1}{4}(1+\frac{1}{n})^2\\
L(f,p)= \sum^n_{r=1}m_rdr=\sum_{r=1}^n \frac{(r-1)^3}{n^3}\frac{1}{n}=\frac{1}{4}(1-\frac{1}{n})^2"


"\int_a^{\bar{b}}f(x)dx = Lim_{n\to \infin} U(f,p)= Lim_{n\to \infin} \frac{1}{4}(1+\frac{1}{n})^2=\frac{1}{4}\\
\int_{\bar{a}}^bf(x)dx = Lim_{n\to \infin} L(f,p)= Lim_{n\to \infin} \frac{1}{4}(1-\frac{1}{n})^2=\frac{1}{4}\\
\int_a^{\bar{b}}f(x)dx=\int_{\bar{a}}^bf(x)dx"

Hence the function is Riemann integrable over [0,1]

"\int_0^1 f(x)dx =\frac{1}{4}"